3L of0 .5 M k 2 Cr 2 O 7 solution have to completely reduced in acidic medium .the number of faraday of electricity required will be

3L of0 .5 M k2Cr2 O7 solution have to completely reduced in acidic medium .the number of faraday of electricity required will be
 

Grade:12

1 Answers

Sandeep Bhamoo
5 years ago
N factor for K2Cr2O7 is = 3 hence electrons required per molecule of it need 3 electron now since there are 3/2 mole of it hence current required is = 3×3/2 = 9/2 faraday current. Please do approve my answr if correct